OpenStudy (anonymous):

The hypotenuse of an isosceles right triangle is 8 inches. Find the length of the legs.

OpenStudy (anonymous):

Are the legs of equal length?

OpenStudy (anonymous):

Since the triangle is an isosceles the legs are equal length. Therefore the Pythagorean theorem can be written \[x^{2}+x^{2} = 8^{2}\] Solve for x.

OpenStudy (lgbasallote):

since it's isosceles..the legs are equal..so a=b a^2 + b^2 = c^2 is the pythagorean theorem since a= b...it will become a^2 + a^2 = c^2 2a^2 = c^2 a^2 = (c^2)/2 a = sqrt(c^2)/sqrt 2 so... a = c/sqrt2 a = 8/sqrt 2 rationalize... a = (8sqrt2)/2 a = 4sqrt2 = b